# Juru: Legal Brazilian Large Language Model from Reputable Sources

### v1

@@ Abstract
The high computational cost associated with pretraining large language models limits their research. Two strategies have emerged to address this issue: domain specialization and pretraining with high-quality data. To explore these strategies, we specialized the Sabiá-2 Small model with 1.9 billion unique tokens from reputable Brazilian legal sources and conducted few-shot evaluations on legal and general knowledge exams. Our model, Juru, demonstrates the benefits of domain specialization with a reduced amount of pretraining data. However, this specialization comes at the expense of degrading performance in other knowledge areas within the same language. This study contributes to the growing body of scientific evidence showing that pretraining data selection may enhance the performance of large language models, enabling the exploration of these models at a lower cost.
@@ 1 Introduction
Large Language Models (LLMs) are commonly trained on vast amounts of general-purpose data sourced from internet snapshots, such as the CommonCrawl [1, 2, 3] and C4 [4, 3]. Scaling laws [5] support this paradigm, suggesting more capable models with increased data and training parameters. As a result, LLMs developed through this method demonstrate the ability to address tasks across multiple languages and domains [6].
The significant computational resources required for pretraining with such massive datasets pose a challenge for research, limiting the exploration of their capabilities. However, it is possible to observe competitive results on a smaller scale through the application of continued pretraining. In the literature, the selection of domain-specific data [7, 8, 9] from reputable sources [10, 11, 12] stand out.
The large general-purpose pretraining of LLMs promotes general factual knowledge and language comprehension. Therefore, continued pretraining on the same scale and scope might not be necessary to improve LLMs’ performance. By employing domain-specific data from reputable sources, we postulate that a LLM initially trained on general-purpose data can improve its performance in a specific domain with minimal additional data, although this may come at the expense of performance degradation in other domains.
To test this hypothesis, we specialized the Sabiá-2 Small [13] model for the Brazilian legal domain, utilizing $1.9$ billion unique tokens sourced mainly from academic studies from reputable sources, resulting in the Juru model. We evaluated the performance of our model on benchmarks of multiple-choice questions in Brazilian law and general knowledge. Our objective is to examine the impact of incorporating only domain-specific knowledge during the model training process.
Across the proposed pretraining, our model demonstrated enhanced performance on the law benchmark, while exhibiting a decrease in performance on the general knowledge benchmark. This contributes to the growing body of scientific evidence supporting the effectiveness of specialized pretraining for improving the performance of generalist models on specific domains. Such specialization may facilitate their research by reducing the computational cost associated with training.
2 Related works
@@
Language and domain specialization have become an effective practice for ensuring that LLMs bring tangible benefits in real-world applications. Recent works have shown that continued pretrained in language-specific corpora, particularly in Portuguese [9, 13, 15], enhances significantly the performance of downstream tasks without the need to increase the parameter count.
@@
SaulLM-7B [7] is the first LLM pretrained to tackle the nuances of legal languages and follow instructions to support conversation interactions. Starting from Mistral-7B [17], the model is pretrained on a large corpus of English-centered documents of legal domain. To the best of our knowledge, Juru is the first LLM pretrained for the Brazilian legal domain.
@@ 3.1 Pretraining Data
A total of $13,281$ documents were extracted from the academic paper databases. Nearly $99\%$ of these documents are in Portable Document Format (PDF). Hence, we used PyMuPDF [^2] https://github.com/pymupdf/PyMuPDF for extracting text from these documents.
Texts from $14,743$ federal laws were extracted from the LexML database. Additionally, the dataset from the study by [18] was added to the pretraining dataset, consisting of $32,535$ legal documents distributed between decisions and judgments from the Brazilian Supreme Federal Court.
@@
Datasets Tokens
Academic papers $1,806,023,152$
@@
Total $1,999,721,118$
Table 1: Distribution of Byte-Pair Encoding (BPE) [20] tokens for the pretraining dataset, using the Sabiá-2 Small tokenizer.
Considering the tokenizer of Sabiá-2 Small, the model employed in our pretraining, we gathered $1.9$ billion tokens, as shown in Table 1. Each document was divided into sequences of $4,096$ tokens.
@@ 3.2 Juru Model
The Juru [^3] Juru is the Tupi name for the largest Brazilian parrot, in homage to the Sabiá model, which served as the main foundation for this research. model extends the causal language modeling pretraining of a Sabiá-2 Small model specialized in Portuguese, without instruction fine-tuning. We employed the t5x and seqio frameworks [21] for the proposed pretraining.
The pretraining hyperparameters are based on the Sabiá model [9]. We employed the AdaFactor optimizer [22] without factorization, with a first-order momentum of $0.9$ and a second-order momentum of $1-k^{-0.8}$, where $k$ represents the current training step. Additionally, we applied dynamic weight decay by $lr^{2}$, with $lr$ as the current learning rate value, and global norm clipping of $1.0$.
In addition to the loss associated with the causal language modeling task, we utilized an auxiliary loss of $10^{-4}\log^{2}\left(\sum_{i}e^{z_{i}}\right)$, where $z$ denotes the logits of the prediction. This auxiliary loss helps dampen the values of the pretraining loss. Initially, the learning rate is set at 00 and gradually increases during a linear warm-up period of $100$ training steps until it reaches $0.001$. Following this warm-up phase, the learning rate remains constant.
The model was pretrained for causal language modeling on a cluster of TPUs v2-128. Each batch consisted of $512$ sequences, with $4,096$ tokens in each sequence. The pretraining process ran for $2,800$ training steps, equivalent to $2.94$ epochs, processing a total of $5.88$ billion tokens. This process took $44.22$ hours, achieving a Model FLOPs Utilization (MFU) of $55.5\%$ [23], excluding self-attention operations.
@@ 4 Evaluation
We aim to evaluate the dynamics of learning domain-specific knowledge while also retaining the knowledge previously learned in the pretraining phrase. Therefore, we conducted two main evaluations of the Juru model, focusing on Brazilian law and general knowledge with Brazilian standardized multiple-choice exams in Portuguese.
4.1 Law
For evaluating the Juru model, we utilized questions from two Bar Association Exams (OAB) of $2023$ and the National Student Performance Exam (ENADE) of $2022$. The OAB exam is obligatory for all law graduates aspiring to practice law in Brazil. It comprises two phases; the initial phase encompasses $80$ multiple-choice questions, while the second phase consists of $4$ essay questions and the creation of a procedural document. ENADE aims to assess the performance of undergraduate students in the subjects outlined for their respective courses. This exam includes $35$ multiple-choice questions, and $5$ essay questions.
The evaluation of law knowledge was conducted with a single example for few-shot learning, using accuracy as the evaluation metric. We included only multiple-choice questions from the $2023$ OAB exams and the ENADE in law, comprising a total set of $195$ questions, as illustrated in Table 2.
Benchmark Task Metric Size
OAB-2023 Multiple-choice (4) Accuracy 160
ENADE-2022-Law Multiple-choice (5) Accuracy 35
Table 2: Configuration of law knowledge benchmarks.
4.2 General Knowledge
Pires et al. [9] observed a performance degradation in English benchmarks after the specialization of an English-centric LLM for Portuguese. Analog to their study, the following question arises: does this trend persist when training a Portuguese LLM for the Brazilian legal domain?
Thus, we explored the impact of domain specialization on the general knowledge of the Juru model with the following Brazilian Portuguese multiple-choice exams [13]:
  • Accounting (CFCES);
  • Medicine (Revalidation and residency exams for USP and UNICAMP);
  • Computer science (POSCOMP);
  • Undergraduate admission (USP, UNICAMP, and ENEM);
  • College education evaluation (ENADE 2022 and 2023).
@@ 4.2 General Knowledge
Following [13], we ignore all the questions that require image understanding. Some exams have captions describing the image contents. In those cases, we incorporate them into the exact position the images would appear.
@@ 5 Results
The benchmarks tracked the performance progression of the pretrained model at each checkpoint. Figure 1 illustrates the study’s hypothesis: as the LLM specializes, domain-specific performance increases while general knowledge performance declines. Upon analyzing the curves, we observed a notable increase in the models’ scores for law benchmarks compared to general knowledge. The law curve peaked at a maximum accuracy score of $72.0\%$ at $3.78$B tokens, whereas the general knowledge curve reached its maximum score of $68.8\%$ at the Sabiá-2 Small model.
Figure 1: Accuracy on law and general knowledge benchmarks with respect to the number of training tokens related to legal Brazilian texts. One epoch is $1.9$ billion tokens. The first points (0B) correspond to the Sabiá-2 Small model.
The optimal version of Juru was reached at $3.78$B tokens, displaying the highest accuracy for the law benchmark. The next subsection compares the performance of Sabiá-2 Small with the model pretrained with $3.78$B tokens for both law and general knowledge benchmarks.
@@ 5.1 Results per Benchmark
In Table 3, we observe an increase of approximately $6$ points in accuracy when comparing Sabiá-2 Small with the model pretrained on $3.78$B tokens on the law benchmarks. All law benchmarks on $3.78$B tokens show higher scores. Notably, the ENADE-2022-Law benchmark presented the highest improvement of the pretrained model, with an increase of about $11$ points compared to Sabiá-2 Small.
@@
Benchmark Sabiá-2 Small Juru
OAB-2023 $60.0\%$ $62.5\%$
ENADE-2022-Law $70.4\%$ $81.5\%$
Mean $65.2\%$ $72.0\%$
Table 3: Comparison of Sabiá-2 Small against Juru on law benchmarks.
Figure 2 presents the comparison between Juru and Sabiá-2 Small across $53$ areas of knowledge, solely for the ENADE of 2022 and 2023. With our pretraining, we observe a decrease in performance for $48$ exams. The Speech Therapy exam presents the highest decline with a decrease of $14$ points in accuracy. Even with a decline in performance for most exams, we observed a sharp increase for the medicine exam with $13$ points in accuracy.
Figure 2: Comparison of Sabiá-2 Small against Juru on the general knowledge benchmark with ENADE 2022 and 2023.
@@ 5.2 Results Analysis
The results demonstrate the dynamics of learning domain-specific knowledge for the specialization of the Juru model. Specialization in Brazilian law yielded a significant increase of $6$ points compared to Sabiá-2 Small in the law benchmark, despite the relatively small dataset of only $1.9$ billion unique tokens.
@@
In line with the hypothesis of this study, specialization has led to a decline in the model’s capacity to solve tasks beyond its specialized domain. Despite the shared language among evaluated tasks, specialization in only one knowledge domain has resulted in decreased performance in other areas. Clear correlations between types of knowledge areas and degradation were not evident.
@@
We posit that further pretraining of the Juru model is likely to increase the disparity between law and general knowledge curves. Consequently, while performance in law-related tasks may increase, there would likely be a significant decrease in performance in specific knowledge domains that require different skill sets.
5.3 Data Contamination

### v2

@@ Abstract
The high compute cost associated with pretraining large language models limits their research. Two strategies have emerged to address this issue: domain specialization and pretraining with high-quality data. To explore these strategies, we specialized the Mistral-7B model with 1.9 billion unique tokens from reputable Brazilian legal sources and conducted few-shot evaluations on legal and general knowledge test suites. Our model, Juru, demonstrates the benefits of domain specialization by achieving improved performance on legal benchmarks, even with a reduced amount of pretraining data. However, this domain specialization through continued pretraining comes at the cost of increased forgetting in unrelated domains, as evidenced by performance degradation on general knowledge test suites in both Portuguese and English. This study contributes to the growing body of scientific evidence showing that pretraining data selection may enhance the performance of large language models, enabling the exploration of these models at a lower cost. Juru is publicly available at https://huggingface.co/roseval/Juru-7B.
Keywords:
Domain Specialization Continued Pretraining Legal Language Models Low-Resource Pretraining
@@ 1 Introduction
Large Language Models (LLMs) are commonly trained on vast amounts of general-purpose data sourced from internet snapshots, such as the CommonCrawl [3, 4, 32] and C4 [32, 34]. Scaling laws [14] support this paradigm, suggesting more capable models with increased data and trainable parameters. As a result, LLMs developed through this method demonstrate the ability to address tasks across multiple languages and domains [20].
The significant computational resources required for pretraining with such massive datasets pose a challenge for research, limiting the exploration of their capabilities. However, it is possible to observe competitive results on a smaller scale through the application of continued pretraining. In the literature, the selection of domain-specific data [8, 9, 22] from reputable sources [12, 18, 21] stands out.
The large general-purpose pretraining of LLMs promotes general factual knowledge and language comprehension. Therefore, continued pretraining on the same scale and scope might not be necessary to improve LLMs’ performance. By employing domain-specific data from reputable sources, we postulate that an LLM initially trained on general-purpose data can improve its performance in a specific domain with minimal additional data, although this may come at the expense of performance degradation in other domains.
To test this hypothesis, we specialized the Mistral-7B [15] model for the Brazilian legal domain, using 1.91.91.9 billion unique tokens sourced mainly from academic studies from reputable sources. The resulting model, Juru, was evaluated on multiple-choice test suites covering both Brazilian legal knowledge and general knowledge domains. Our objective is to assess the impact of incorporating exclusively domain-specific data during continued pretraining on the model’s performance.
Across the proposed pretraining, our model demonstrated enhanced performance on the legal test suite, while exhibiting a decrease in performance on the general knowledge test suites. In particular, a more pronounced decrease was observed on the English test suite compared to the Portuguese one, suggesting that previously acquired knowledge is better retained when it shares greater similarity with the target domain of specialization. These results contribute to the growing body of scientific evidence supporting the effectiveness of specialized pretraining for improving the performance of generalist models on specific domains. Such specialization may facilitate their research by reducing the compute cost associated with training. Additionally, the resulting model, Juru, is publicly released to support further research on domain specialization through continued pretraining. To the best of our knowledge, Juru is the first LLM pretrained for the Brazilian legal domain.
2 Related Work
@@
Language and domain specialization have become effective strategies for ensuring that LLMs bring tangible benefits in real-world applications. Recent works have shown that continued pretrained in language-specific corpora, particularly in Portuguese [1, 17, 22], enhances significantly the performance of downstream tasks without the need to increase the parameter count.
@@
SaulLM-7B [8] is the first LLM pretrained to tackle the nuances of legal languages and follow instructions to support conversation interactions. Starting from Mistral-7B [15], the model was pretrained on 30 billion tokens of an English legal dataset. In a subsequent study, Colombo et al. [7] scaled this approach to a Mixture-of-Experts architecture. The authors continually pretrained Mixtral-54B and Mixtral-141B on 540 billion tokens of English legal texts, producing SaulLM-54B and SaulLM-141B. While these larger models achieve higher performance on most in-domain tasks, their increased size correlates with lower performance on certain legal tasks.
To the best of our knowledge, Juru is the first LLM pretrained for the Brazilian legal domain. Our experiments show that, relative to its base model, Juru achieves superior performance on in-domain tasks. However, its performance deteriorates on tasks beyond the Brazilian legal domain.
@@ 3.1 Pretraining Data
A total of 13,278 documents were extracted from academic paper databases. Nearly $\%9$ of these documents are in Portable Document Format (PDF). Hence, we used Marker [^2] https://github.com/VikParuchuri/marker for extracting text from these documents.
Texts from 14,743 federal laws were extracted from the LexML database. Additionally, the dataset from the study by Sakiyama et al. [27] was added to the pretraining dataset, consisting of 32,535 legal documents distributed between decisions and judgments from the Brazilian Supreme Federal Court.
@@
Table 1: Distribution of Byte-Pair Encoding (BPE) [29] tokens for the pretraining dataset, using the Mistral-7B tokenizer.
Dataset Tokens
Academic papers 1,805,059,922
@@
Total 1,998,757,888
Considering the tokenizer of Mistral-7B, the model employed in our pretraining, we gathered 1.91.91.9 billion tokens, as shown in Table 1. Each document was divided into sequences of 4,096 tokens.
@@ 3.2 Juru Model
The Juru model extends the causal language modeling pretraining of a Mistral-7B model, without instruction finetuning. We employed the t5x and seqio frameworks [26] for the proposed pretraining.
Although more recent models [10, 11] of comparable size have demonstrated superior performance compared to Mistral-7B, we chose Mistral-7B due to its earlier release date. This decision aims to reduce the risk of data contamination in our test suites.
For the pretraining hyperparameters, we employed the AdaFactor optimizer [30] without factorization, with a first-order momentum of 0.90.90.9 and a second-order momentum of $1-k^{-0.8}$, where $k$ represents the current training step. Additionally, we applied dynamic weight decay by $lr^{2}$, with $lr$ as the current learning rate value, and global norm clipping of 1.01.01.0.
In addition to the loss associated with the causal language modeling task, we used an auxiliary loss of $10^{-4}\log^{2}\left(\sum_{i}e^{z_{i}}\right)$, where $z$ denotes the logits of the prediction. This auxiliary loss helps dampen the values of the pretraining loss. The learning rate is initially set at 0 and gradually increases during a linear warm-up period of 250250250 training steps until it reaches 0.0010.0010.001. Following this warm-up phase, the learning rate remains constant.
[…]
@@ 4 Evaluation
We investigate how continued pretraining through domain specialization influences both knowledge learning and forgetting. Hence, we evaluate the specialized Juru model on three standardized multiple-choice test suites: Brazilian legal knowledge, representing the target domain of our specialization; Portuguese general knowledge, which share the same language as the target domain but do not assess legal knowledge; and English general knowledge, whose distribution is closer to the original English-centric training data of Mistral-7B.
For all test suites, we adopt accuracy as the primary evaluation metric. We group all editions of the same exam administered within a given year into a single evaluation group. When aggregating multiple exams, we report the mean accuracy. Lastly, we employ few-shot learning with three examples during our evaluation, as the model was trained solely for text completion and did not undergo any instruction finetuning.
4.1 Brazilian Legal Knowledge
For evaluating the Juru model, we used questions from six Bar Association Exams (OAB) of 202420242024 and 202320232023, and two National Judiciary Exams (ENAM) of 202420242024. The OAB exam is mandatory for all law graduates aspiring to practice law in Brazil. It comprises two phases; the initial phase encompasses 808080 multiple-choice questions, while the second phase consists of 444 essay questions and the creation of a procedural document. ENAM was introduced in 2023 to unify the selection process for entry into the Brazilian judiciary. As of its implementation, only candidates who have passed the ENAM within the last two years are eligible to apply for judicial positions in Brazil. This exam includes a single phase with 808080 multiple-choice questions. We included only multiple-choice questions from the 202320232023 and 202420242024 OAB, and 202420242024 ENAM exams, comprising a total set of 638638638 questions, as illustrated in Table 2.
Table 2: Configuration of Brazilian legal knowledge test suite.
Benchmark Domain Exams Task Size
OAB-2023 Legal 3 Multiple-choice (4) 240
OAB-2024 Legal 3 Multiple-choice (4) 238
ENAM-2024 Legal 2 Multiple-choice (5) 160
4.2 Portuguese General Knowledge
Pires et al. [22] observed a performance degradation on English benchmarks following the specialization of an English-centric LLM for Portuguese. These results raise the following question: Is catastrophic forgetting limited to cross-linguistic domain specialization, or does specializing in a specific domain within a single language also lead to forgetting of general knowledge in that language?
To address this question, we examined the impact of specializing the Juru model in the Brazilian legal domain on its performance in Portuguese general knowledge tasks, using the multiple-choice exams listed in Table 3, comprising a total amount of 2,123 questions.
Table 3: Configuration of Portuguese general knowledge test suite.
Benchmark Domain Exams Task Size
ENEM-2024 General 1 Multiple-choice (5) 179
BLUEX-2024 General 2 Multiple-choice (5) 172
CPNU-2024 General 7 Multiple-choice (5) 320
BNDES-2024 General 12 Multiple-choice (5) 420
REVALIDA-2024 Medical 2 Multiple-choice (4) 187
MREX-2024 Medical 6 Multiple-choice (4) 452
CFCEQ-2024 Accounting 12 Multiple-choice (5) 294
CFCES-2024 Accounting 2 Multiple-choice (4) 99
@@ 4.2 Portuguese General Knowledge
Following Almeida et al. [1], we ignore all the questions that require image understanding. Some exams have captions describing the image contents. In those cases, we incorporate them into the exact position the images would appear.
4.3 English General Knowledge
Unlike the Sabiá [22] model, which underwent extensive continued pretraining, the Juru model was trained for fewer steps due to the smaller size of its legal specialization dataset. While Pires et al. [22] observed a loss of English knowledge in their model after extended training on Portuguese data, their experiments involved a significantly longer training regime and a more diverse dataset. This raises another question: Can shorter, domain-specific pretraining mitigate or prevent catastrophic forgetting of previously learned knowledge?
To investigate this, we evaluated the Juru model on the MMLU benchmark [13], a widely used test suite for assessing general knowledge and reasoning abilities in English [15, 20, 31]. While MMLU includes a broad range of subjects, such as STEM, humanities, law, and medicine, we restricted our evaluation to the college-level and high school-level subsets, amounting to 4,369 questions, as outlined in Table 4.
Table 4: Configuration of English general knowledge test suite.
Benchmark Domain Exams Task Size
MMLU-College General 6 Multiple-choice (4) 719
MMLU-High School General 14 Multiple-choice (4) 3,650
@@ 5 Results
The proposed test suites were used to track the model’s performance across checkpoints during continued pretraining. Figure 1 presents the accuracy as a function of the number of tokens processed during continued pretraining. Each epoch comprises 1.91.91.9 billion tokens from Brazilian legal texts. The 0B point corresponds to the base Mistral-7B model, while the checkpoint at 7.1B tokens represents the specialized Juru model.
Figure 1: Accuracy on the Brazilian legal knowledge, English general knowledge, and Portuguese general knowledge test suites during continued pretraining.
[…]
Overall, the optimal version of Juru was obtained at 7.1B tokens, corresponding to the highest accuracy on the legal knowledge test suite. All subsequent analyses are based on this checkpoint, which we refer to as Juru. The next subsection compares the performance of this final model to the base Mistral-7B model across both legal and general knowledge test suites.
@@ 5.1 Results per Benchmark
Table 5 shows that Juru outperforms the base Mistral-7B model across all legal benchmarks, with an overall gain of approximately 4.7% in average accuracy. The most notable improvement is observed in the ENAM-2024 benchmark, with a gain of 5.6%. This result demonstrates the effectiveness of domain specialization through continued pretraining in enhancing performance on legal knowledge tasks.
Table 5: Comparison of Mistral-7B against Juru on Brazilian legal knowledge benchmarks.
@@
Benchmark Mistral-7B Juru
OAB-2023 48.3% 54.5%
OAB-2024 49.6% 52.0%
ENAM-2024 31.2% 36.8%
Mean (8 Exams) 44.5% 49.2%
In contrast, Table 6 indicates a decline in performance on the English general knowledge benchmarks after domain specialization. While Mistral-7B achieves higher scores on both the college and high school subsets of MMLU, Juru underperforms across the board, resulting in a 3.6% decrease in the overall average accuracy.
Table 6: Comparison of Mistral-7B against Juru on English general knowledge benchmarks.
Accuracy
Benchmark Mistral-7B Juru
MMLU-College 54.2% 49.4%
MMLU-High School 64.8% 61.6%
Mean (20 Exams) 61.6% 58.0%
A similar trend is observed for the Portuguese general knowledge benchmarks in Table 7. Juru shows lower accuracy in 5 out of 8 individual benchmarks, leading to a drop of 2.4% in the average accuracy across all exams.
Table 7: Comparison of Mistral-7B against Juru on Portuguese general knowledge benchmarks.
Accuracy
Benchmark Mistral-7B Juru
ENEM-2024 62.0% 61.4%
BLUEX-2024 60.7% 58.4%
CPNU-2024 56.8% 56.8%
BNDES-2024 59.1% 56.0%
REVALIDA-2024 50.7% 51.7%
MREX-2024 42.2% 36.7%
CFCEQ-2024 40.7% 38.0%
CFCES-2024 40.3% 41.3%
Mean (44 Exams) 50.5% 48.1%
Figure 2 presents the comparison between Juru and Mistral-7B across 13 areas of knowledge from the BNDES-2024 benchmark. Overall, Juru underperforms the base model in 8 out of the 13 exams. The most significant decline is observed in the IT Support, with a drop of approximately 11.4%. In contrast, Juru maintains the same accuracy as Mistral-7B in IT Cybersecurity, Data Science, and Accounting, and shows gains in Architecture and Economics. Despite these specific improvements, the overall trend highlights a decrease in performance across general knowledge areas following domain-specific continued pretraining.
Figure 2: Comparison of Juru against Mistral-7B on BNDES-2024 exams.
@@ 5.2 Results Analysis
[…]
@@
In line with the hypothesis of this study, specialization has led to a decline in the model’s capacity to solve tasks beyond its specialized domain. Despite the shared language among evaluated tasks on the legal and Portuguese general test suites, specialization in only one knowledge domain has resulted in decreased performance in other domain areas. Clear correlations between types of knowledge areas and degradation were not evident, as shown in Figure 2. Furthermore, when comparing the decline in performance across the Portuguese and English general knowledge test suites, the Juru model exhibited a greater degree of forgetting on the English suite. This pattern suggests that knowledge forgetting is influenced by the similarity between the evaluated and specialized domains: the less similar the previously learned knowledge is to the target domain, the greater the degree of forgetting.
@@
We posit that further pretraining of the Juru model is likely to increase the disparity between legal and general knowledge curves. Consequently, while performance in legal-related tasks may increase, there would likely be a significant decrease in performance in specific knowledge domains that require different skill sets.
6 Limitations